OpenStudy (anonymous):

The difference between 1/2 o a number and 1/5 of the same number is 24. Find the number

OpenStudy (zzr0ck3r):

the difference in two numbers is |a-b| or the larger minus the smaller. so you have the difference of two numbers one is (1/2) of some number and the other is (1/5) of some number 1/2>1/5 so 1/2x-1/5x=24

OpenStudy (anonymous):

set \[\frac{ 1 }{ 2 }x-\frac{1}{5}x=24\]

OpenStudy (anonymous):

Multiply everythnig by 10 to get rid of the fractions:\[5x-2x=240\] and solve for x

OpenStudy (anonymous):

Oh...i get it now ^^ Thanks guys. (:

OpenStudy (zzr0ck3r):

np